OpenStudy (anonymous):

how many six letter words can be formed from the letters A,B,C,D,E,F, if each letter can only be used once?

OpenStudy (anonymous):

720. If you want a detailed explanation, just read on If you want the maximum number of possible combinations of those letters, you'll want to progressively multiply them together. It works like this: Lets say the number of letters equals N and the total number of answers is X by progressively multiplying we get: X=N*(N-1)*(N-2)*(N-3)*(N-4)*...*1 so for N=6 (i.e. A,B,C,D,E&F) X=6*(6-1)*(6-2)*(6-3)*(6-4)*(6-5) X=6*5*4*3*2*1 X=720 If you have a scientific calculator, you can easily calculate this by entering the number of letters (i.e. 6) and pressing the (x!) button